OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find 2f(1) + 3g

OpenStudy (bibby):

let's start with the first half. what is f(1)? f(x) = x^2 + 1 f(1) = 1^2 + 1 = ?

OpenStudy (anonymous):

2

OpenStudy (bibby):

so what is 2f(1)?

OpenStudy (anonymous):

4

OpenStudy (bibby):

I think you didn't copy the question right. 3g doesn't mean anything. g of what?